# Synesthesia of Machines (SoM)-Enhanced ISAC Precoding for Vehicular Networks with Double Dynamics

### v1

@@ II-C Sensing Model
We model the target’s trajectory as a unidirectional curve with an orientation offset. $\left[\theta_{\text{s}}^{(n)},d_{\text{s}}^{(n)}\right]$ denotes the target’s position in polar domain. The target response matrix at the $m$-th subcarrier $\bm{G}^{(n)}_{\text{m}}=\alpha^{(n)}_{\text{m}}\bm{a}_{\text{r}}(\theta_{\text{s}}^{(n)})\bm{a}_{\text{t}}^{\text{H}}(\theta_{\text{s}}^{(n)})$, where $\theta_{\text{s}}^{(n)}$ is the azimuth of the target relative to the BS and $\alpha_{\text{m}}^{(n)}$ denotes the complex reflection coefficient incorporating the round-trip path loss, the radar cross-section and the Doppler of the target. Accordingly the received echo becomes
$\bm{y}_{\text{r,m}}^{(n,l)}=\bm{G}_{\text{m}}^{(n)}\bm{x}_{\text{m}}^{(n,l)}+\bm{z}_{\text{m}}^{(n,l)},\vspace{-0.2cm}$ (9)
with $\bm{z}_{\text{m}}^{(n,l)}\sim\mathcal{CN}(\bm{0},\sigma_{\text{s}}^{2}\bm{I}_{N_{\text{r}}})$ being the ISAC receiver noise. Note that BS also receives clutter reflected by objects unrelated to the target in the environment. We consider whitening the received noise to ensure that the noise obeys the complex Gaussian distribution.
$\mathcal{J}_{\text{m}}(\theta_{\text{s}}^{(n)})=\frac{2\|\alpha_{m}\|^{2}(\text{Tr}(\dot{\bm{A}}^{\text{H}}(\theta_{\text{s}}^{(n)})\dot{\bm{A}}(\theta_{\text{s}}^{(n)})\bm{R}_{\text{x,m}}^{(n)})\text{Tr}(\bm{A}^{\text{H}}(\theta_{\text{s}}^{(n)})\bm{A}(\theta_{\text{s}}^{(n)})\bm{R}_{\text{x,m}}^{(n)})-\|\text{Tr}(\dot{\bm{A}}^{\text{H}}(\theta_{\text{s}}^{(n)})\bm{A}(\theta_{\text{s}}^{(n)})\bm{R}_{\text{x,m}}^{(n)})\|^{2})}{\sigma_{\text{s}}^{2}\text{Tr}(\bm{A}^{\text{H}}(\theta_{\text{s}}^{(n)})\bm{A}(\theta_{\text{s}}^{(n)})\bm{R}_{\text{x,m}}^{(n)})}.$ (10)
The sensing accuracy is characterized by Fisher information in Eq. (10), with $\bm{R}_{\text{x,m}}^{(n)}=\frac{1}{L}\mathbb{E}\left[\sum_{l=1}^{L}\bm{x}_{\text{m}}^{(n,l)}\bm{x}_{\text{m}}^{(n,l),\text{H}}\right]=\frac{P_{\text{t}}}{N_{\text{t}}UM}\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB,m}}^{(n)}{\bm{F}_{\text{BB,m}}^{(n)}}^{\text{H}}{\bm{F}_{\text{RF}}^{(n)}}^{\text{H}}$ being the covariance matrix of $\bm{x}_{\text{m}}$, and $\bm{A}(\theta_{\text{s}})=\bm{a}_{\text{r}}(\theta_{\text{s}})\bm{a}_{\text{t}}^{\text{H}}(\theta_{\text{s}})$ with $\dot{\bm{A}}(\theta_{\text{s}})=\frac{\partial\bm{A}(\theta_{\text{s}})}{\partial\theta_{\text{s}}}$. The Fisher information for $\theta_{\text{s}}$ at $t$ is the summation of all the Fisher information across subcarriers, as analyzed in [27]. Consequently, the Cramér-Rao lower bound (CRLB) of the target’s azimuth is the expectation of the inverse of Fisher information given $\bm{y}_{\text{r}}$.
$\text{CRLB}(\theta_{\text{s}}^{(n)})\!=\!\mathbb{E}\left[\frac{1}{\mathcal{J}(\theta_{\text{s}}^{(n)})}\right]\!=\!\mathbb{E}\left[\frac{1}{\sum_{m=1}^{M}\mathcal{J}_{\text{m}}(\theta_{\text{s}}^{(n)})}\right].\vspace{-0.2cm}$ (11)
@@ III-B Alternating-based Optimization
The goal of the hybrid precoding in this doubly-dynamic scenario is to maximize ISAC utility throughout the trajectory, considering the power budget and constant-modulus constraint of the analog precoder. The utility is defined as the weighted sum of the averaged spectrum efficiency $\overline{\mathcal{R}}$ and the averaged Fisher information for angle estimation $\overline{\mathcal{J}}=\frac{1}{T}\sum_{n=1}^{T}\sum_{m=1}^{M}\mathcal{J}_{\text{m}}(\theta_{\text{s}}^{n})$. Both indicators are normalized by the ideal performance boundaries ¯$\overline{\mathcal{R}}^{*}=\frac{1}{T}\sum_{n=1}^{T}{\mathcal{R}^{(n)}}^{*}$ and $\overline{\mathcal{J}}^{*}=\frac{1}{T}\sum_{n=1}^{T}{\mathcal{J}^{(n)}}^{*}$ respectively, with ${\mathcal{R}^{(n)}}^{*}$ and ${\mathcal{J}^{(n)}}^{*}$ being the performance boundaries in the $t$-th subframe, obtained by individual maximization of SE without sensing constraints, and individual minimization of CRLB without communication requirements, respectively. Consequently, the problem can be formulated as
@@
s.t. $~{}\sum_{m=1}^{M}\|\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB,m}}^{(n)}\|_{\text{F}}^{2}=1,$ (12a)
@@
given the perfect $\overline{\bm{H}}$ and perfect $\bm{G}_{\text{m}}$ in each subframe. $\psi$ denotes the weighting coefficient between the sensing and communication performances. (12a) and (12b) denote the power constraint and the constant-modulus constraint respectively. The problem (12) is non-convex due to the complex objective function and constant-modulus constraints. Note that the time-averaged objective can be maximized when the utility at each single subframe is maximized. We first transform the sequential optimization on a subframe-by-subframe basis as
@@
s.t. $~{}\sum_{m=1}^{M}\|\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB,m}}^{(n)}\|_{\text{F}}^{2}=1,$ (13a)
$~{}|[\bm{F}_{\text{RF}}^{(n)}]_{i,j}|=1.$ (13b)
To tackle the coupled variables and non-convex constraints, we adopt an alternating-based optimization procedure. We first optimize the equivalent fully-digital precoding matrices $\bm{F}_{\text{m}}^{(n)}=\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB,m}}^{(n)}$ for ISAC utility maximization, then optimize $\bm{F}_{\text{RF}}^{(n)}$ and $\bm{F}_{\text{BB,m}}^{(n)}$ by decomposing $\bm{F}_{\text{m}}^{(n)}$ in an alternating manner. Specifically, in each alternating iteration, the phases in $\bm{F}_{\text{RF}}^{(n)}$ are updated with fixed $\bm{F}_{\text{BB,m}}^{(n)}$, followed by updating $\bm{F}_{\text{BB,m}}^{(n)}$ with $\bm{F}_{\text{RF}}^{(n)}$ held constant.
@@
$\underset{\bm{F}_{\text{m}}}{\text{max}}$ $~{}\psi\frac{\mathcal{R}}{\mathcal{R}^{*}}+(1-\psi)\frac{\mathcal{J}}{\mathcal{J}^{*}}$ (14)
s.t. $~{}\sum_{m=1}^{M}\|\bm{F}_{\text{m}}\|_{\text{F}}^{2}=1.$ (14a)
Let $\mathcal{J}_{\text{m}}(\theta_{\text{s}})\geqslant\zeta$ and $\gamma_{\text{m,u}}\geqslant\tau$, $\forall m,u$. Denote $\bm{W}_{\text{m,u}}=\bm{f}_{\text{m,u}}\bm{f}_{\text{m,u}}^{\text{H}}$, $\bm{W}_{\text{m}}=\sum_{u=1}^{U}\bm{W}_{\text{m,u}}$ and $\bm{Q}_{\text{m,k,u}}=\bm{h}_{\text{m,k,u}}^{\text{H}}\bm{h}_{\text{m,k,u}}$. Then the problem (14) can be transformed into
$\underset{\bm{W}_{\text{m,u}}}{\text{max}}$ $~{}\omega$ (15)
s.t. $~{}\bm{W}_{\text{m,u}}\succeq\bm{0},~{}~{}\text{rank}(\bm{W}_{\text{m,u}})=1,~{}$ (15a)
@@
$~{}\left[\begin{matrix}\text{Tr}(\dot{\bm{A}}^{\text{H}}\dot{\bm{A}}\bm{W}_{\text{m}})-\zeta&\text{Tr}(\dot{\bm{A}}^{\text{H}}\bm{A}\bm{W}_{\text{m}})\\ \text{Tr}(\bm{A}^{\text{H}}\dot{\bm{A}}\bm{W}_{\text{m}})&\text{Tr}(\bm{A}^{\text{H}}\bm{A}\bm{W}_{\text{m}})\end{matrix}\right]\geq\bm{0},$ (15c)
$~{}\text{Tr}(\bm{Q}_{\text{m,m,u}}\bm{W}_{\text{m,u}})-\tau\sum_{i=1,i\neq u}^{U}\!\text{Tr}(\bm{Q}_{\text{m,m,u}}\bm{W}_{\text{m,i}})$
$~{}-\tau\sum_{k=1,k\neq m}\bm{Q}_{\text{m,k,u}}\bm{W}_{\text{k,u}}\leqslant\tau\sigma_{\text{c}}^{2},$ (15d)
[…]
where (15c) represents for the sensing constraint by applying Schur-complement theorem, (15b) represents for the SINR constraint, and $\omega$ serves as the lower bound for the ISAC utility per subframe. Subsequently, we employ the classical SDR technique by eliminating the rank-1 constraints in (15a). This problem is then transformed into a standard semi-definite programming (SDP) and can be solved by the off-the-shelf toolbox like CVX. As a result, the rank-1 constraint is restored as elaborated in [10] and the equivalent fully-digital precoding matrix $\bm{F}_{m}$ is obtained.
@@
Upon obtaining $\bm{F}_{\text{m}}$, the remaining task is to optimize $\bm{F}_{\text{BB,m}}$ and $\bm{F}_{\text{RF}}$ to approximate $\bm{F}_{\text{m}}$ as
$\underset{\bm{F}_{\text{BB,m}},\bm{F}_{\text{RF}}}{\text{min}}$ $~{}\sum_{m=1}^{M}\|\bm{F}_{\text{m}}-\bm{F}_{\text{RF}}\bm{F}_{\text{BB,m}}\|_{\text{F}}^{2}$ (16)
@@ III-B1 Optimizing $\bm{F}_{\text{RF}}$
$\sum_{m=1}^{M}\|\bm{F}_{\text{m}}-\bm{F}_{\text{RF}}\bm{F}_{\text{BB,m}}\|_{\text{F}}^{2}$ $=\!\sum_{m=1}^{M}\|(\bm{F}_{\text{m}}\bm{F}_{\text{BB,m}}^{-1}-\bm{F}_{\text{RF}})\bm{F}_{\text{BB,m}}\|_{\text{F}}^{2}$ (17)
$\leqslant\!\sum_{m=1}^{M}\|\bm{F}_{\text{m}}\bm{F}_{\text{BB,m}}^{-1}-\bm{F}_{\text{RF}}\|_{\text{F}}^{2}\|\bm{F}_{\text{BB,m}}\|_{\text{F}}^{2},$
@@
$\underset{\bm{F}_{\text{RF}}}{\text{min}}$ $~{}\sum_{m=1}^{M}\|\bm{F}_{\text{m}}\bm{F}_{\text{BB,m}}^{-1}-\bm{F}_{\text{RF}}\|_{\text{F}}^{2}\|\bm{F}_{\text{BB,m}}\|_{\text{F}}^{2}$ (18)
s.t. $~{}|[\bm{F}_{\text{RF}}]_{i,j}|=1.$ (18a)
@@
$\bm{F}_{\text{RF}}=\text{exp}\{j\angle(\sum_{m=1}^{M}\|\bm{F}_{\text{BB,m}}\|_{\text{F}}^{2}\bm{F}_{\text{m}}\bm{F}_{\text{BB,m}}^{-1})\}.\vspace{-0.2cm}$ (19)

### v2

@@ II-C Sensing Model
We model the target’s trajectory as a unidirectional curve with an orientation offset. $\left[\theta_{\text{s}}^{(n)},d_{\text{s}}^{(n)}\right]$ denotes the target’s position in polar domain. The target response matrix at the $m$-th subcarrier $\bm{G}^{(n)}_{m}=\alpha^{(n)}_{m}\bm{a}_{\text{r}}(\theta_{\text{s}}^{(n)})\bm{a}_{\text{t}}^{\text{H}}(\theta_{\text{s}}^{(n)})$, where $\theta_{\text{s}}^{(n)}$ is the azimuth of the target relative to the BS and $\alpha_{m}^{(n)}$ denotes the complex reflection coefficient incorporating the round-trip path loss, the radar cross-section and the Doppler of the target. Accordingly the received echo becomes
$\bm{y}_{\text{r},m}^{(n,l)}=\bm{G}_{m}^{(n)}\bm{x}_{m}^{(n,l)}+\bm{z}_{m}^{(n,l)},\vspace{-0.2cm}$ (9)
with $\bm{z}_{m}^{(n,l)}\sim\mathcal{CN}(\bm{0},\sigma_{\text{s}}^{2}\bm{I}_{N_{\text{r}}})$ being the ISAC receiver noise. Note that BS also receives clutter reflected by objects unrelated to the target in the environment. We consider whitening the received noise to ensure that the noise obeys the complex Gaussian distribution.
$\mathcal{J}_{m}(\theta_{\text{s}}^{(n)})=\frac{2\|\alpha_{m}\|^{2}(\text{Tr}(\dot{\bm{A}}^{\text{H}}(\theta_{\text{s}}^{(n)})\dot{\bm{A}}(\theta_{\text{s}}^{(n)})\bm{R}_{x,m}^{(n)})\text{Tr}(\bm{A}^{\text{H}}(\theta_{\text{s}}^{(n)})\bm{A}(\theta_{\text{s}}^{(n)})\bm{R}_{x,m}^{(n)})-\|\text{Tr}(\dot{\bm{A}}^{\text{H}}(\theta_{\text{s}}^{(n)})\bm{A}(\theta_{\text{s}}^{(n)})\bm{R}_{x,m}^{(n)})\|^{2})}{\sigma_{\text{s}}^{2}\text{Tr}(\bm{A}^{\text{H}}(\theta_{\text{s}}^{(n)})\bm{A}(\theta_{\text{s}}^{(n)})\bm{R}_{x,m}^{(n)})}.$ (10)
The sensing accuracy is characterized by Fisher information in Eq. (10), with $\bm{R}_{x,m}^{(n)}=\frac{1}{L}\mathbb{E}\left[\sum_{l=1}^{L}\bm{x}_{m}^{(n,l)}\bm{x}_{m}^{(n,l),\text{H}}\right]=\frac{P_{\text{t}}}{N_{\text{t}}UM}\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB},m}^{(n)}{\bm{F}_{\text{BB},m}^{(n)}}^{\text{H}}{\bm{F}_{\text{RF}}^{(n)}}^{\text{H}}$ being the covariance matrix of $\bm{x}_{m}$, and $\bm{A}(\theta_{\text{s}})=\bm{a}_{\text{r}}(\theta_{\text{s}})\bm{a}_{\text{t}}^{\text{H}}(\theta_{\text{s}})$ with $\dot{\bm{A}}(\theta_{\text{s}})=\frac{\partial\bm{A}(\theta_{\text{s}})}{\partial\theta_{\text{s}}}$. The Fisher information for $\theta_{\text{s}}$ at $t$ is the summation of all the Fisher information across subcarriers, as analyzed in [31]. Consequently, the Cramér-Rao lower bound (CRLB) of the target’s azimuth is the expectation of the inverse of Fisher information given $\bm{y}_{\text{r}}$.
$\text{CRLB}(\theta_{\text{s}}^{(n)})\!=\!\mathbb{E}\left[\frac{1}{\mathcal{J}(\theta_{\text{s}}^{(n)})}\right]\!=\!\mathbb{E}\left[\frac{1}{\sum_{m=1}^{M}\mathcal{J}_{m}(\theta_{\text{s}}^{(n)})}\right].\vspace{-0.2cm}$ (11)
@@ III-B Alternating-based Optimization
The goal of the hybrid precoding in this doubly-dynamic scenario is to maximize ISAC utility throughout the trajectory, considering the power budget and constant-modulus constraint of the analog precoder. The utility is defined as the weighted sum of the averaged spectrum efficiency $\overline{\mathcal{R}}$ and the averaged Fisher information for angle estimation $\overline{\mathcal{J}}=\frac{1}{T}\sum_{n=1}^{T}\sum_{m=1}^{M}\mathcal{J}_{m}(\theta_{\text{s}}^{n})$. Both indicators are normalized by the ideal performance boundaries ¯$\overline{\mathcal{R}}^{*}=\frac{1}{T}\sum_{n=1}^{T}{\mathcal{R}^{(n)}}^{*}$ and $\overline{\mathcal{J}}^{*}=\frac{1}{T}\sum_{n=1}^{T}{\mathcal{J}^{(n)}}^{*}$ respectively, with ${\mathcal{R}^{(n)}}^{*}$ and ${\mathcal{J}^{(n)}}^{*}$ being the performance boundaries in the $t$-th subframe, obtained by individual maximization of SE without sensing constraints, and individual minimization of CRLB without communication requirements, respectively. Consequently, the problem can be formulated as
@@
s.t. $~{}\sum_{m=1}^{M}\|\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB},m}^{(n)}\|_{\text{F}}^{2}\leqslant 1,$ (12a)
@@
given the perfect $\overline{\bm{H}}$ and perfect $\bm{G}_{m}$ in each subframe. $\psi$ denotes the weighting coefficient between the sensing and communication performances. (12a) and (12b) denote the power constraint and the constant-modulus constraint respectively. The problem (12) is non-convex due to the complex objective function and constant-modulus constraints. Note that the time-averaged objective can be maximized when the utility at each single subframe is maximized. We first transform the sequential optimization on a subframe-by-subframe basis as
@@
s.t. $~{}\sum_{m=1}^{M}\|\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB},m}^{(n)}\|_{\text{F}}^{2}\leqslant 1,$ (13a)
$~{}|\bm{F}_{\text{RF}}^{(n)}[i,j]|=1.$ (13b)
To tackle the coupled variables and non-convex constraints, we adopt an alternating-based optimization procedure. We first optimize the equivalent fully-digital precoding matrices $\bm{F}_{m}^{(n)}=\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB},m}^{(n)}$ for ISAC utility maximization, then optimize $\bm{F}_{\text{RF}}^{(n)}$ and $\bm{F}_{\text{BB},m}^{(n)}$ by decomposing $\bm{F}_{m}^{(n)}$ in an alternating manner. Specifically, in each alternating iteration, the phases in $\bm{F}_{\text{RF}}^{(n)}$ are updated with fixed $\bm{F}_{\text{BB},m}^{(n)}$, followed by updating $\bm{F}_{\text{BB},m}^{(n)}$ with $\bm{F}_{\text{RF}}^{(n)}$ held constant.
@@
$\underset{\bm{F}_{m}}{\text{max}}$ $~{}\psi\frac{\mathcal{R}}{\mathcal{R}^{*}}+(1-\psi)\frac{\mathcal{J}}{\mathcal{J}^{*}}$ (14)
s.t. $~{}\sum_{m=1}^{M}\|\bm{F}_{m}\|_{\text{F}}^{2}\leqslant 1.\vspace{-0.2cm}$ (14a)
Denote $\bm{W}_{m,u}=\bm{f}_{m,u}\bm{f}_{m,u}^{\text{H}}$, $\bm{W}_{m}=\sum_{u=1}^{U}\bm{W}_{m,u}$ and $\bm{Q}_{m,k,u}=\bm{h}_{m,k,u}^{\text{H}}\bm{h}_{m,k,u}$. Then the problem (14) can be transformed into
$\underset{\bm{W}_{m,u},\zeta,\tau}{\text{max}}$ $~{}\omega$ (15)
s.t. $~{}\bm{W}_{m,u}\succeq\bm{0},~{}~{}\text{rank}(\bm{W}_{m,u})=1,~{}$ (15a)
@@
$~{}\left[\begin{matrix}\text{Tr}(\dot{\bm{A}}^{\text{H}}\dot{\bm{A}}\bm{W}_{m})-\zeta&\text{Tr}(\dot{\bm{A}}^{\text{H}}\bm{A}\bm{W}_{m})\\ \text{Tr}(\bm{A}^{\text{H}}\dot{\bm{A}}\bm{W}_{m})&\text{Tr}(\bm{A}^{\text{H}}\bm{A}\bm{W}_{m})\end{matrix}\right]\geq\bm{0},$ (15c)
$~{}\text{Tr}(\bm{Q}_{m,m,u}\bm{W}_{m,u})-\tau\sum_{i=1,i\neq u}^{U}\!\text{Tr}(\bm{Q}_{m,m,u}\bm{W}_{m,i})$
$~{}-\tau\sum_{k=1,k\neq m}\bm{Q}_{m,k,u}\bm{W}_{k,u}\geqslant\tau\sigma_{\text{c}}^{2},$ (15d)
$~{}\sum_{m=1}^{M}\text{Tr}(\bm{W}_{m})\leqslant 1.$ (15e)
where (15b) represents for the SINR constraint equivalent to $\gamma_{m,u}\geqslant\tau$, since $|\bm{h}_{m,m,u}\bm{f}_{m,u}|^{2}=\text{Tr}(\bm{h}_{m,m,u}\bm{W}_{m,u}\bm{h}_{m,m,u}^{\text{H}})$. (15c) represents for the sensing constraint equivalent to $\mathcal{J}_{m}(\theta_{\text{s}})\geqslant\zeta$, by applying Schur complement theorem. $\omega$ serves as the lower bound for the ISAC utility per subframe. Subsequently, we employ the classical SDR technique by eliminating the rank-1 constraints in (15a). This problem is then transformed into a standard semi-definite programming (SDP) and can be solved by the off-the-shelf toolbox like CVX. As a result, the rank-1 constraint is restored as elaborated in [10] and the equivalent fully-digital precoding matrix $\bm{F}_{m}$ can be obtained.
Proposition 1.
There exists a global optimum for the relaxed problem satisfying the rank-1 constraint, which ensures the tightness of the SDR can be guaranteed.
Proof.
See Appendix A. ∎
@@
Upon obtaining $\bm{F}_{m}$, the remaining task is to optimize $\bm{F}_{\text{BB},m}$ and $\bm{F}_{\text{RF}}$ to approximate $\bm{F}_{m}$ as
$\underset{\bm{F}_{\text{BB},m},\bm{F}_{\text{RF}}}{\text{min}}$ $~{}\sum_{m=1}^{M}\|\bm{F}_{m}-\bm{F}_{\text{RF}}\bm{F}_{\text{BB},m}\|_{\text{F}}^{2}$ (16)
@@ III-B1 Optimizing $\bm{F}_{\text{RF}}$
$\sum_{m=1}^{M}\|\bm{F}_{m}-\bm{F}_{\text{RF}}\bm{F}_{\text{BB},m}\|_{\text{F}}^{2}$ $=\!\sum_{m=1}^{M}\|(\bm{F}_{m}\bm{F}_{\text{BB},m}^{-1}-\bm{F}_{\text{RF}})\bm{F}_{\text{BB},m}\|_{\text{F}}^{2}$ (17)
$\leqslant\!\sum_{m=1}^{M}\|\bm{F}_{m}\bm{F}_{\text{BB},m}^{-1}-\bm{F}_{\text{RF}}\|_{\text{F}}^{2}\|\bm{F}_{\text{BB},m}\|_{\text{F}}^{2},$
@@
$\underset{\bm{F}_{\text{RF}}}{\text{min}}$ $~{}\sum_{m=1}^{M}\|\bm{F}_{m}\bm{F}_{\text{BB},m}^{-1}-\bm{F}_{\text{RF}}\|_{\text{F}}^{2}\|\bm{F}_{\text{BB},m}\|_{\text{F}}^{2}$ (18)
s.t. $~{}|\bm{F}_{\text{RF}}[i,j]|=1.$ (18a)
@@
$\bm{F}_{\text{RF}}=\text{exp}\{j\angle(\sum_{m=1}^{M}\|\bm{F}_{\text{BB},m}\|_{\text{F}}^{2}\bm{F}_{m}\bm{F}_{\text{BB},m}^{-1})\}.\vspace{-0.2cm}$ (19)